OpenStudy (-welp-):

"Simplify. (4x2 – 3x + 2) – (2x2 + 5x – 7)" How do I solve this?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Subtracting polynomials is like subtracting real numbers. How do you subtract two real numbers a and b in algebra? $$\Huge a-b = ?$$

OpenStudy (skullpatrol):

To subtract b you add the opposite of b: $$\Huge a-b = a+(-b)$$ for example $$2-4 = 2 + (-4)$$

OpenStudy (skullpatrol):

To subtract a polynomial, you add the opposite of that polynomial and then simplify.

OpenStudy (skullpatrol):